# SQL-Factory: A Multi-Agent Framework for High-Quality and Large-Scale SQL Generation

### v1

@@ 5.1.3. Baseline
SQL Generation Baselines. To compare the quality and diversity of generated SQL queries, we consider both rule-based and learning-based generation methods. We include SQLSmith (Seltenreich, 2016), a grammar-driven query generator that produces random SQLs over a given schema, and two representative learning-based methods: LearnedSQLGen (Zhang et al., 2022), which learns to generate queries via reinforcement learning, and OmniSQL (Li et al., 2025a), which leverages large language models to synthesize large-scale SQL datasets. Additionally, we exclude template-based methods (e.g., TPC-DS-style generators) from our comparison, as they rely heavily on predefined query templates and are not suitable for generating massive and structurally diverse SQL workloads.
Text-to-SQL Baselines. For evaluating the effectiveness of the generated data on downstream Text-to-SQL tasks, we fine-tune two base models: Qwen2.5-Coder-1.5B and Qwen2.5-Coder-3B, using SQL-NL pairs derived from the SQL queries synthesized by SQL-Factory. The NL annotations are obtained via a post-processing step, as described in Section 5.4. We compare the performance of these fine-tuned models with several model of similar scale, including Qwen2.5-Coder-7B (Hui et al., 2024), Qwen2.5-7B (Yang et al., 2024), Deepseek-Coder-6.7B (Guo et al., 2024), Codes-1B, and Codes-3B (Li et al., 2024). This setup allows us to examine whether augmenting training with generated queries generated by SQL-Factory can improve model performance, even when using relatively smaller model backbones.
@@ 5.1.4. Implementation Details
We implement all components of the SQL-Factory framework with a combination of large language models, retrieval-based tools, and external SQL execution engines. Each module is configured to balance quality, efficiency, and ease of integration.
@@
Expansion Agent. For the Expansion Agent, we utilize Qwen2.5-Coder-14B (Hui et al., 2024), a locally deployed open-source language model optimized for code and SQL-related instruction following. It is served using vLLM to support efficient batch decoding at scale, enabling rapid query expansion. It is also use a temperature of 0.8 for diversity.
Decision Agents. The Management Agent, Table Selection Agent, Seed Selection Agent and Critical Agent also reused Qwen2.5-Coder-14B to reduce deployment overhead. Here we use a lower temperature of 0.3, aiming for more deterministic and stable reasoning behavior without sacrificing overall coherence.
Similarity and Execution Tools. For hybrid similarity calculation, we use a combination of three tools: a BERT-large (Devlin et al., 2019) encoder for dense embeddings, the rapidfuzz library for token-level overlap, and sqlglot library for AST-level structural comparison. For executability validation, the Execution Tool runs each SQL query against a SQLite engine and uses the EXPLAIN command to check for successful parsing and plan generation.
5.2. Main Results
@@
5.2.1. Generation Stopping Criteria
The generation process proceeds in alternating rounds of generation and expansion, as described in Section 4.1. To determine when to terminate for each benchmark, we monitor two key metrics: (1) the expansion ratio, defined as the proportion of all queries in the SQL Pool that originate from the Expansion Team, and (2) the average hybrid similarity, which reflects query redundancy in the overall pool.
@@ 5.2.1. Generation Stopping Criteria
Figure 5. Trend chart of the proportion of Expansion and overall similarity with the change in the total number of queries.
As illustrated in Figure 5, the generation process is initialized by the Generation Team, which produces a small number of diverse seed queries. After this warm-up phase, the Expansion Team begins to take over, resulting in a rise in both the expansion ratio and average similarity across the SQL Pool. This indicates that a growing number of queries are being synthesized based on previously explored patterns.
5.2.2. Statistics of Generated Queries
Despite the varying generation needs, SQL-Factory maintains high efficiency. The total cost across all four benchmarks remains under \$200. This efficiency is achieved by coupling a small number of high-quality queries generated via GPT-4o with scalable expansions performed using Qwen2.5-Coder model. Together, these results demonstrate that SQL-Factory is capable of producing high-quality SQL workloads at scale, while maintaining both schema adaptivity and cost-effectiveness.
@@ 5.2.3. Comparative Analysis on Query Diversity
We conduct a comparative study to assess the query diversity achieved by SQL-Factory against three representative SQL generation methods: SQLSmith (Seltenreich, 2016), a grammar-based random query generator; LearnedSQLGen (Zhang et al., 2022), which employs a reinforcement learning framework guided by cost-based feedback to generate SQL queries through a finite-state abstraction; and OmniSQL (Li et al., 2025a), a LLM-based framework that leverages prompt engineering to generate millions of executable SQL queries.
Hybrid Similarity. We use our proposed hybrid similarity metric as described in Section 4.1.2 to measure the average pairwise similarity between queries. This metric integrates token-level, structural-level, and embedding-level similarity signals, providing a robust measure of redundancy in query content and structure. Lower scores indicate higher diversity. As shown in Table 2, SQL-Factory achieves a balanced similarity score of 0.637, significantly lower than LearnedSQLGen (0.863) and OmniSQL (0.688), suggesting that it avoids over-repetition while maintaining meaningful structure. Besides, the low similarity score for SQLSmith stems largely from its use of excessive syntactic randomness. It frequently inserts arbitrary user-defined functions, opaque predicates, and irrelevant clauses that contribute to lexical variance but not to meaningful semantic or structural diversity.
Vendi Score. The Vendi Score (Friedman and Dieng, 2023) is a recently proposed metric for evaluating diversity based on the exponential of the entropy of the eigenvalues of a similarity matrix computed from the data. Unlike traditional diversity metrics, it does not rely on predefined labels or reference distributions, making it especially suitable for open-ended generation tasks. We utilize Vendi Score to assess the diversity of our generated datasets with two embedding backbones: SimCSE (Gao et al., 2021), a general-purpose sentence encoder trained via contrastive learning, and SQL-Encoder (Pourreza et al., 2024), a specialized model pretrained for SQL embeddings. Higher Vendi scores indicate greater overall diversity in the embedding space. As shown in Table 2, SQL-Factory achieves the highest Vendi Score under both encoders, clearly outperforming other baselines. In contrast, SQLSmith demonstrates low similarity but also fails to produce coherent queries, which is reflected in its relatively low Vendi Score. LearnedSQLGen, while maintaining executable outputs, suffers from high redundancy due to its pattern imitation and optimization-driven training. Besides, OmniSQL is limited by its monolithic generation strategy and lack of internal evaluation during synthesis.
Table 2. Comparison with Existing Queries Generation Methods on TPC-DS.
@@
5.3. Case Study of Generated SQL Queries
@@
Figure 6. Evaluation of Generated Query Distribution.
We conduct a detailed case study to evaluate the quality of the SQL queries synthesized by SQL-Factory. Inspired by analysis methods used in (Zhang et al., 2022), we evaluate 40,000 queries generated for the IMDB benchmark and report their distribution across several key dimensions. These dimensions include join depth, predicate usage, subquery nesting, aggregate usage, and token length, each of which reflects different aspects of query expressiveness and syntactic variety. The results are summarized in Figure 6.
Figure 6(a) illustrates the distribution of join operations per query. Notably, 74% of all queries involve at least one join, and 35% contain two or more join. This indicates that SQL-Factory effectively generates multi-table queries, showcasing its capacity for relational reasoning and schema-aware generation. In particular, 8.3% of the queries feature three or more joins, highlighting the system’s ability to construct complex cross-table logic. Figure 6(b) analyzes the number of predicates in queries. More than 60% of the generated queries contain two or more predicates, with 11.8% having four or more. This confirms that SQL-Factory not only generates structurally rich queries but also embeds fine-grained filtering logic, which is critical for supporting meaningful data retrieval and realistic workload simulation. Besides, Figure 6(c) and Figure 6(d) present the proportions of queries that contain nested subqueries and aggregation operations, respectively. These results demonstrate that a large portion of the generated SQLs utilize advanced SQL constructs, further validating the model’s ability to produce structurally rich and expressive queries. Finally, Figure 6(e) reports the distribution of token lengths across these queries. The x-axis represents the number of tokens per SQL query, while the y-axis shows the corresponding frequency. The distribution reveals a broad range of query lengths, confirming that SQL-Factory does not produce templated or overly uniform queries, but rather supports diverse query formulations at varying levels of complexity.
@@ 5.4.1. Dataset Construction
Instead of generating SQL from NL, which often suffers from semantic mismatch and low exact match accuracy, we adopt a reverse formulation: converting SQL to NL. Recent studies (Wu et al., 2021; Guo et al., 2018; Zhang et al., 2023) have shown that LLMs are more reliable in generating faithful natural language descriptions given SQL, especially in complex multi-table settings. This makes the SQL-to-NL approach a robust and efficient means for constructing high-quality data. To this end, we develop a multi-stage LLM-based pipeline for SQL-to-NL conversion with consistency checking and optional SQL refinement. The steps are as follows:
@@
(3) Semantic Consistency Verification. To further ensure the correctness of the constructed ¡NL, SQL¿ pairs, we introduce an LLM-based verifier to perform semantic consistency checking. This component is critical for filtering out misaligned instances where the generated natural language question does not faithfully reflect the intent or logic of the corresponding SQL query. Specifically, the verifier takes as input the SQL query and its corresponding natural language question, and is prompted to return a binary decision indicating whether the pair is semantically aligned. To improve reasoning quality and reduce false positives, we adopt a Chain-of-Thought (CoT) prompting strategy. Only those ¡NL, SQL¿ pairs that pass this verification step are retained, ensuring high semantic fidelity in the final dataset.
@@ 5.4.2. Model Training
We fine-tune two instruction-tuned backbone models, Qwen2.5-Coder-1.5B and Qwen2.5-Coder-3B, on the constructed dataset of ¡NL, SQL¿ pairs. Both models are pretrained for code and SQL reasoning, making them suitable for Text-to-SQL adaptation. Training is performed using the AdamW optimizer with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\epsilon=1\times 10^{-8}$. We use a peak learning rate of $3\times 10^{-6}$, linearly warmed up over the first 5% of training steps, followed by cosine learning rate decay. Each model is trained for 3 epochs with a batch size of 256 and a maximum sequence length of 4096 tokens. To stabilize training and prevent overfitting, we apply gradient clipping with a threshold of 1.0 and set the weight decay to 0.01. All other settings, including data shuffling and tokenization, follow the default configuration of LLaMA-Factory (Zheng et al., 2024).
@@ 5.4.3. Result and Analysis
To ensure fairness, we adopt the official prompt templates for each model. For codes models, we use the released prompt from the original repository but remove schema linking and value retrieval modules to match our setup. For all Qwen-family and DeepSeek models, we adopt the default instruction prompt used by Qwen2.5-Coder (Hui et al., 2024). Each model is evaluated on two benchmarks: Spider and BIRD, both widely used for assessing Text-to-SQL performance across varying difficulty levels. The evaluation metrics include both executability and execution accuracy.
The results are shown in Table 3. After fine-tuning with the SQL-Factory dataset, both Qwen2.5-Coder-1.5B and Qwen2.5-Coder-3B exhibit substantial performance improvements. On Spider, Qwen2.5-Coder-1.5B+SQL-Factory achieves 97.20% in executability and 76.50% in accuracy, outperforming the DeepSeek-Coder-6.7B and Qwen2.5-7B variants in terms of accuracy. Similarly, Qwen2.5-Coder-3B+SQL-Factory reaches 97.29% executability and 79.69% accuracy, ranking among the strongest models despite using fewer parameters. On BIRD, which contains more complex and compositional queries, fine-tuned models again show superior performance. Qwen2.5-Coder-3B+SQL-Factory achieves 88.07% in executability and 46.09% in accuracy, clearly outperforming all baseline models of similar or larger scale. These improvements are especially pronounced in the “Moderate” and “Challenging” categories, confirming the effectiveness of SQL-Factory in providing structurally rich and semantically aligned training data.
These findings underscore SQL-Factory’s utility in enabling small and mid-sized models to compete with, or even outperform, larger models by leveraging high-quality, diverse, and executable SQL-NL training pairs.
@@ 5.5. Query Clustering Task
Table 4 presents the BetaCV scores across all models and benchmarks. Traditional rule-based methods such as Aouiche, Aligon, and Makiyama rely on handcrafted SQL features or syntax trees, and often fail to capture deep semantic similarities across structurally different but semantically related queries. PreQR improves upon this by leveraging a pretrained language model, but it lacks task-specific supervision and may underfit domain-specific patterns. Meanwhile, untuned BERT models benefit from general semantic priors but are still agnostic to SQL-specific structure and vocabulary.

### v2

@@ 5.1.3. Baseline
SQL Generation Baselines. To compare the quality and diversity of generated SQL queries, we consider both rule-based and learning-based generation methods. We include SQLSmith (Seltenreich, 2016), a grammar-driven query generator that produces random SQLs over a given schema, and two representative learning-based methods: LearnedSQLGen (Zhang et al., 2022), which learns to generate queries via reinforcement learning, and OmniSQL (Li et al., 2025a), which leverages large language models to synthesize large-scale queries. Additionally, we exclude template-based methods (Nambiar and Poess, 2006; Poess and Floyd, 2000; Leis et al., 2018) from our comparison, as they rely heavily on predefined templates and are not suitable for generating massive and structurally diverse SQL workloads.
Text-to-SQL Baselines. For evaluating the effectiveness of the generated data on downstream Text-to-SQL tasks, we fine-tune two base models: Qwen2.5-Coder-1.5B and Qwen2.5-Coder-3B, using SQL-NL pairs derived from the SQL queries synthesized by SQL-Factory. The NL annotations are obtained via a post-processing step, as described in Section 5.3. We compare the performance of these fine-tuned models with several model of similar scale, including Codes-1B, and Codes-3B (Li et al., 2024). This setup allows us to examine whether augmenting training with generated queries generated by SQL-Factory can improve model performance.
@@ 5.1.4. Implementation Details
We utilize different large language models and configurations to implement these agents in SQL-Factory. Each module is configured to balance quality, efficiency, and ease of integration.
@@
Expansion Agent. For the Expansion Agent, we utilize Qwen2.5-Coder-14B (Hui et al., 2024), a locally deployed open-source model optimized for code and SQL-related instruction following. It is served using vLLM to support efficient batch decoding at scale, enabling rapid query expansion. It is also use a temperature of 0.8 for diversity.
Other Agents. The Management Agent, Critical Agent, Table Selection Agent, and Seed Selection Agent also reused Qwen2.5-Coder-14B to reduce deployment overhead. Here we use a lower temperature of 0.3, aiming for more deterministic and stable reasoning behavior without sacrificing overall coherence.
5.2. Quality Analysis of Generated SQL
@@
5.2.1. Overview of Generated Queries
@@
Figure 6. Overview of SQL queries synthesized by SQL-Factory.
Key aspects such as join depth, predicate usage, subquery nesting, aggregate usage, and overall token length provide meaningful insights into the expressiveness and structural diversity of SQL queries. These dimensions serve as proxies for measuring the coverage of different logical constructs and the capability to reflect real-world query distributions. Therefore, we conduct a comprehensive analysis of the SQL queries along these dimensions in Figure 6.
Figure 6(a) illustrates the distribution of join operations per query. Notably, 78.6% of all queries involve at least one join, and 30% contain two or more join. This indicates that SQL-Factory effectively generates multi-table queries, showcasing its capacity for relational reasoning and schema-aware generation. Figure 6(b) analyzes the number of predicates in queries. More than 58% of the generated queries contain two or more predicates, with 9.8% having four or more. This confirms that SQL-Factory not only generates structurally rich queries but also embeds fine-grained filtering logic, which is critical for supporting meaningful data retrieval and realistic workload simulation. Besides, Figure 6(c) and Figure 6(d) present the proportions of queries that contain nested subqueries and aggregation operations, respectively. These results demonstrate that a large portion of the generated SQLs utilize advanced SQL constructs, further validating the model’s ability to produce structurally rich and expressive queries. Finally, Figure 6(e) reports the distribution of token lengths across these queries. The x-axis represents the number of tokens per SQL query, while the y-axis shows the corresponding frequency. The distribution reveals a broad range of query lengths, confirming that SQL-Factory does not produce templated or overly uniform queries, but rather supports diverse query formulations at varying levels of complexity.
5.2.2. Cost Analysis of SQL Generation
@@
Despite the varying generation needs, SQL-Factory maintains high efficiency. The total cost across all four benchmarks remains under \$200. Notably, if the entire workload were generated solely by GPT-4o without the Expansion Team, the API cost would approach \$1000, highlighting the substantial cost savings enabled by our hybrid design. Together, these results demonstrate that SQL-Factory is capable of producing SQL workloads at scale, while maintaining both schema adaptivity and cost-effectiveness.
@@ 5.2.3. Comparative Analysis on Query Diversity
Table 2. Comparison with Existing Query Generation Methods in Terms of Diversity.
@@
We conduct a comparative study to assess the query diversity achieved by SQL-Factory against three representative SQL generation methods: SQLSmith (Seltenreich, 2016), a grammar-based random query generator; LearnedSQLGen (Zhang et al., 2022), which employs a reinforcement learning framework guided by cost-based feedback to generate queries through a finite-state abstraction; and OmniSQL (Li et al., 2025a), a LLM-based framework that leverages prompt engineering to generate millions of executable SQL queries.
Hybrid Similarity. We use our proposed hybrid similarity metric as described in Section 4.1.2 to measure the average pairwise similarity between queries. This metric integrates token-level, structural-level, and embedding-level similarity signals, providing a robust measure of redundancy in query content and structure. Lower scores indicate higher diversity. As shown in Table 2, SQL-Factory achieves a similarity score of 0.637, significantly lower than LearnedSQLGen and OmniSQL, suggesting that it avoids over-repetition while maintaining meaningful structure. Besides, the low similarity score for SQLSmith stems largely from its use of excessive syntactic randomness. It frequently inserts arbitrary user-defined functions, opaque predicates, and irrelevant clauses that contribute to lexical variance but not to meaningful semantic or structural diversity.
Vendi Score. The Vendi Score (Friedman and Dieng, 2023) is a recently proposed metric for evaluating diversity based on the exponential of the entropy of the eigenvalues of a similarity matrix computed from the data. Unlike traditional diversity metrics, it does not rely on predefined labels or reference distributions, making it especially suitable for open-ended generation tasks. We utilize Vendi Score to assess the diversity of our generated datasets with two embedding backbones: SimCSE (Gao et al., 2021), a general-purpose sentence encoder trained via contrastive learning, and SQL-Encoder (Pourreza et al., 2024), a specialized model pretrained for SQL embeddings. Higher Vendi scores indicate greater overall diversity in the embedding space. As shown in Table 2, SQL-Factory achieves the highest Vendi Score under both encoders. In contrast, SQLSmith demonstrates low similarity but also fails to produce coherent queries, which is reflected in its relatively low Vendi Score. LearnedSQLGen suffers from high redundancy due to its pattern imitation and optimization-driven training. Besides, OmniSQL is limited by its monolithic generation strategy.
5.2.4. Schema-Aware Query Allocation Analysis
((a))
((b))
Figure 7. Comparison of query distribution across tables in OmniSQL and SQL-Factory. The x-axis indicates the number of queries per table, and the y-axis shows the proportion of tables within each range.
To assess whether SQL-Factory effectively achieves schema-aware query generation, we compare its table-level query distribution with queries generated by OmniSQL (Li et al., 2025a). As illustrated in Figure 7 (a), OmniSQL exhibits a highly skewed distribution, where a small number of tables dominate the query workload, while most tables receive very few or no queries. This imbalance leads to insufficient schema coverage and hampers the generalizability of downstream models. In contrast, Figure 7 (b) shows that SQL-Factory generates queries with a significantly more balanced distribution. Most tables receive a similar number of queries, typically centered around 400, forming a bell-shaped curve that approximates a normal distribution. This stands in stark contrast to the long-tail distribution observed in OmniSQL. Such uniformity not only reflects SQL-Factory’s schema-aware generation strategy but also promotes broader schema utilization, enabling the synthesis of more representative and robust SQL datasets.
This pattern aligns with the general complexity distribution of tables in real-world databases, which often follows a bell-shaped curve, as most tables exhibit moderate complexity. SQL-Factory’s query allocation reflects this property, assigning more queries to moderately complex tables while avoiding over- or under-representation of extreme cases.
5.2.5. Generation Stopping Criteria
@@
Figure 8. Evolution of Expansion Ratio and Overall Similarity as the Number of Synthesized SQL Queries Increases.
To determine when to terminate for each benchmark, we monitor two key metrics: (1) the expansion ratio, defined as the proportion of all queries in the SQL Pool that originate from the Expansion Team, and (2) the average hybrid similarity, which reflects query redundancy in the overall pool.
As illustrated in Figure 8, the system is initialized by the Generation Team, which produces a small number of diverse seed queries. After this warm-up phase, the Expansion Team begins to take over, resulting in a rise in both the expansion ratio and similarity across the SQL Pool. This indicates that a growing number of queries are being synthesized based on previously explored patterns.
@@ 5.3.1. Dataset Construction
Instead of generating SQL from NL, which often suffers from semantic mismatch and low exact match accuracy, we adopt a reverse formulation: converting SQL to NL. Recent studies (Wu et al., 2021; Guo et al., 2018; Zhang et al., 2023) have shown that LLMs are more reliable in generating faithful natural language descriptions for given SQL. This makes the SQL-to-NL approach a robust and efficient means for constructing high-quality data. To this end, we develop a multi-stage pipeline for SQL-to-NL conversion with consistency checking and optional SQL refinement. The steps are as follows:
@@
(3) Semantic Consistency Verification. To further ensure the correctness of the constructed ¡NL, SQL¿ pairs, we introduce an LLM-based verifier to perform semantic consistency checking. This component is critical for filtering out misaligned instances where the generated natural language question does not faithfully reflect the intent or logic of the corresponding SQL query. Specifically, the verifier takes as input the SQL query and its corresponding natural language question, and is prompted to return a binary decision indicating whether the pair is semantically aligned. To improve reasoning quality, we adopt a Chain-of-Thought (CoT) prompting strategy. Only those ¡NL, SQL¿ pairs that pass this verification step are retained, ensuring high semantic fidelity in the final dataset.
@@ 5.3.2. Model Training
We fine-tune two instruction-tuned backbone models, Qwen2.5-Coder-1.5B and Qwen2.5-Coder-3B, on the constructed dataset of ¡NL, SQL¿ pairs. Both models are pretrained for code and SQL reasoning, making them suitable for Text-to-SQL adaptation. Training is performed using the AdamW optimizer with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\epsilon=1\times 10^{-8}$. We use a peak learning rate of $3\times 10^{-6}$, linearly warmed up over the first 5% of training steps, followed by cosine learning rate decay. Each model is trained for 3 epochs with a batch size of 256 and a maximum sequence length of 4096 tokens. To stabilize training and prevent overfitting, we apply gradient clipping with a threshold of 1.0 and set the weight decay to 0.01. All other settings follow the default configuration of LLaMA-Factory (Zheng et al., 2024).
@@ 5.3.3. Result and Analysis
To ensure fairness, we adopt the official prompt templates for each model. For codes models, we use the released prompt from the original repository but remove schema linking and value retrieval modules to match our setup. For all Qwen-family models, we adopt the default instruction prompt used by Qwen2.5-Coder (Hui et al., 2024). Each model is evaluated on two benchmarks: Spider and BIRD and the evaluation metrics include both executability and execution accuracy.
The results are shown in Table 3. After fine-tuning with the SQL-Factory dataset, both Qwen2.5-Coder-1.5B and Qwen2.5-Coder-3B exhibit substantial performance improvements. On Spider, Qwen2.5-Coder-1.5B+SQL-Factory achieves an executability of 97.20% and accuracy of 76.50%, while Qwen2.5-Coder-3B+SQL-Factory reaches 97.29% executability and 79.69% accuracy. Both outperform their corresponding base models by a large margin, demonstrate that SQL-Factory effectively boosts model performance even without increasing model size. On BIRD, which contains more complex and compositional queries, fine-tuned models again show superior performance. Qwen2.5-Coder-3B+SQL-Factory achieves 88.07% in executability and 46.09% in accuracy, clearly outperforming all baseline models of similar or larger scale. These improvements are especially pronounced in the “Moderate” and “Challenging” categories, confirming the effectiveness of SQL-Factory in providing structurally rich and semantically aligned training data.
Overall, these findings show that with access to SQL-Factory, small and mid-sized models can achieve strong performance in Text-to-SQL tasks. This highlights SQL-Factory’s ability to deliver diverse and executable training examples that bridge the performance gap without requiring massive model sizes.
@@ 5.4. Query Clustering Task
Table 4 presents the BetaCV scores across all models and benchmarks. Traditional rule-based methods such as Aouiche, Aligon, and Makiyama rely on handcrafted SQL features or syntax trees, and often fail to capture deep semantic similarities across structurally different but semantically related queries. PreQR improves upon this by leveraging a pretrained model, but it lacks task-specific supervision and may underfit domain-specific patterns. Meanwhile, untuned BERT models benefit from general semantic priors but are still agnostic to SQL-specific structure and vocabulary.